A steel wire of length 64 cm weighs 5 g. If it is stretched by a force of 8 N, what would be the speed
of, transverse wave passing on it ?
 

Grade:12

1 Answers

Arun
6 years ago
mass = 5 g, length l = 64 cm ∴ mass per unit length = m = 5/64 g/cm ∴ Tension, T = 8N = 8 * 105 dyne V = √(T /m) = √(8 * 105 * 64) / 5 = 3200 cm/s = 32 m/s